# Revisit What You See: Disclose Language Prior in Vision Tokens for Efficient Guided Decoding of LVLMs

### v1

@@ C.1 Case studies for vision token analysis
river $-14.74$ $0.00$ $-6.16$ $0.21$ $-3.61$ $2.70$ $-0.36$ $69.75$ $-14.58$ $0.00$ $-5.02$ $0.66$
mountain $-14.07$ $0.00$ $-5.48$ $0.42$ $-6.47$ $0.15$ $-3.22$ $3.99$ $-15.02$ $0.00$ $-5.46$ $0.43$
person $-13.39$ $0.00$ $-4.80$ $0.82$ $-12.27$ $0.00$ $-9.01$ $0.01$ $-9.86$ $0.01$ $-0.30$ $73.81$
sky $-8.66$ $0.02$ $-0.07$ $93.44$ $-6.76$ $0.12$ $-3.51$ $2.99$ $-12.82$ $0.00$ $-3.26$ $3.83$
tree $-12.06$ $0.00$ $-3.48$ $3.09$ $-11.71$ $0.00$ $-8.46$ $0.02$ $-12.47$ $0.00$ $-2.92$ $5.42$
cloud $-13.88$ $0.00$ $-5.30$ $0.50$ $-9.52$ $0.01$ $-6.26$ $0.19$ $-12.46$ $0.00$ $-2.91$ $5.47$
sea $-13.80$ $0.00$ $-5.21$ $0.55$ $-6.82$ $0.11$ $-3.57$ $2.81$ $-13.24$ $0.00$ $-3.68$ $2.52$
grass $-15.41$ $0.00$ $-6.82$ $0.11$ $-7.19$ $0.08$ $-3.94$ $1.95$ $-13.41$ $0.00$ $-3.86$ $2.11$
rock $-13.35$ $0.00$ $-4.77$ $0.85$ $-4.95$ $0.71$ $-1.70$ $18.33$ $-12.41$ $0.00$ $-2.85$ $5.77$
@@
1 painting $-2.52$ $8.08$ $-1.27$ $27.99$ $-3.79$ $39.15$
2 rock $-2.60$ $7.42$ $-2.28$ $10.22$ $-4.88$ $13.11$
3 mountain $-2.14$ $11.76$ $-3.03$ $4.83$ $-5.17$ $9.82$
4 person $-1.89$ $15.10$ $-3.59$ $2.75$ $-5.48$ $7.19$
5 landscape $-2.27$ $10.30$ $-3.42$ $3.27$ $-5.70$ $5.82$
6 d $-3.45$ $3.16$ $-2.29$ $10.14$ $-5.74$ $5.55$
7 p $-3.91$ $2.01$ $-2.45$ $8.60$ $-6.36$ $3.00$
8 scene $-3.80$ $2.23$ $-2.70$ $6.75$ $-6.50$ $2.60$
9 small $-2.83$ $5.91$ $-3.70$ $2.48$ $-6.52$ $2.54$
10 beautiful $-3.67$ $2.54$ $-3.03$ $4.84$ $-6.70$ $2.13$
11 large $-2.93$ $5.34$ $-3.78$ $2.28$ $-6.71$ $2.11$
12 scen $-3.57$ $2.81$ $-3.17$ $4.19$ $-6.74$ $2.04$
13 hill $-4.01$ $1.82$ $-2.83$ $5.91$ $-6.84$ $1.86$
14 chair $-2.85$ $5.78$ $-4.38$ $1.26$ $-7.23$ $1.26$
15 bow $-3.87$ $2.09$ $-3.68$ $2.53$ $-7.54$ $0.92$
16 second $-3.80$ $2.23$ $-4.52$ $1.09$ $-8.33$ $0.42$
17 smaller $-3.02$ $4.90$ $-6.04$ $0.24$ $-9.06$ $0.20$
18 boat $-3.00$ $4.98$ $-6.34$ $0.18$ $-9.34$ $0.15$
19 distant $-4.14$ $1.59$ $-5.43$ $0.44$ $-9.57$ $0.12$
@@ C.2 Without vocabulary subset case study
C.3 Additional qualitative examples

### v2

@@ C.1 Case studies for vision token analysis
[…]
mountain 14.07-14.07 $0.00$ 5.48-5.48 $0.42$ 6.47-6.47 $0.15$ 3.22-3.22 $3.99$ 15.02-15.02 $0.00$ 5.46-5.46 $0.43$
person 13.39-13.39 $0.00$ 4.80-4.80 $0.82$ 12.27-12.27 $0.00$ 9.01-9.01 $9.86 0.010.01 \cellcolorgray$$0.30 \cellcolorgray!20$$73.81$
[…]
tree 12.06-12.06 $0.00$ 3.48-3.48 $3.09$ 11.71-11.71 $0.00$ 8.46-8.46 $0.02$ 12.47-12.47 $0.00$ 2.92-2.92 $5.42$
cloud 13.88-13.88 $0.00$ 5.30-5.30 $0.50$ 9.52-9.52 $0.01$ 6.26-6.26 $0.19$ 12.46-12.46 $0.00$ 2.91-2.91 $5.47$
sea 13.80-13.80 $0.00$ 5.21-5.21 $0.55$ 6.82-6.82 $0.11$ 3.57-3.57 $2.81$ 13.24-13.24 $0.00$ 3.68-3.68 $2.52$
grass 15.41-15.41 $0.00$ 6.82-6.82 $0.11$ 7.19-7.19 $0.08$ 3.94-3.94 $1.95$ 13.41-13.41 $0.00$ 3.86-3.86 $2.11$
rock 13.35-13.35 $0.00$ 4.77-4.77 $0.85$ 4.95-4.95 $0.71$ 1.70-1.70 $18.33$ 12.41-12.41 $0.00$ 2.85-2.85 $5.77$
@@
1 painting 2.52-2.52 $8.08$ 1.27-1.27 $27.99$ 3.79-3.79 $39.15$
2 rock 2.60-2.60 $7.42$ 2.28-2.28 $10.22$ 4.88-4.88 $13.11$
3 mountain 2.14-2.14 $11.76$ 3.03-3.03 $4.83$ 5.17-5.17 $9.82$
4 person 1.89-1.89 $15.10$ 3.59-3.59 $2.75$ 5.48-5.48 $7.19$
5 landscape 2.27-2.27 $10.30$ 3.42-3.42 $3.27$ 5.70-5.70 $5.82$
6 d 3.45-3.45 $3.16$ 2.29-2.29 $10.14$ 5.74-5.74 $5.55$
7 p 3.91-3.91 $2.01$ 2.45-2.45 $8.60$ 6.36-6.36 $3.00$
8 scene 3.80-3.80 $2.23$ 2.70-2.70 $6.75$ 6.50-6.50 $2.60$
9 small 2.83-2.83 $5.91$ 3.70-3.70 $2.48$ 6.52-6.52 $2.54$
10 beautiful 3.67-3.67 $2.54$ 3.03-3.03 $4.84$ 6.70-6.70 $2.13$
11 large 2.93-2.93 $5.34$ 3.78-3.78 $2.28$ 6.71-6.71 $2.11$
12 scen 3.57-3.57 $2.81$ 3.17-3.17 $4.19$ 6.74-6.74 $2.04$
13 hill 4.01-4.01 $1.82$ 2.83-2.83 $5.91$ 6.84-6.84 $1.86$
14 chair 2.85-2.85 $5.78$ 4.38-4.38 $1.26$ 7.23-7.23 $1.26$
15 bow 3.87-3.87 $2.09$ 3.68-3.68 $2.53$ 7.54-7.54 $0.92$
16 second 3.80-3.80 $2.23$ 4.52-4.52 $1.09$ 8.33-8.33 $0.42$
17 smaller 3.02-3.02 $4.90$ 6.04-6.04 $0.24$ 9.06-9.06 $0.20$
18 boat 3.00-3.00 $4.98$ 6.34-6.34 $0.18$ 9.34-9.34 $0.15$
19 distant 4.14-4.14 $1.59$ 5.43-5.43 $0.44$ 9.57-9.57 $0.12$
@@ C.2 Without vocabulary subset case study
C.3 Qualitative examples
To illustrate the qualitative improvements, we present examples from the LLaVA-Bench-In-The-Wild dataset, which consists of diverse real-world images requiring complex reasoning beyond conventional benchmarks. As shown in Figure 9, the input image is a cartoon illustration requiring contextual and conceptual understanding. While baseline methods such as vanilla greedy decoding fail to capture the main theme and generate overly generic or literal descriptions, or even getting stuck in repetitive phrasing. In contrast, ReVisiT successfully identifies the underlying concept: about neural network layers.